OpenStudy (anonymous):

find three points for this equation -x-y=1

OpenStudy (anonymous):

if you set x = 1,2,3 then solve for y you will have your three coordinates i will do the frist one for you :) i.e : \[x=1\]\[-1-y=1\]\[-y =1 + 1\]\[y= -1(1+1)=-2\] since we got y = -2 now we have both our y and x in our first coordinate x=1 y=-2 (1,-2)